Solve the following system of non-linear equations.
y=x^2-2x-33
y=x+7

Answer by ikleyn(52778)

.
In given equations, left sides are identical --- hence, right sides are equal

    x^2 - 2x - 33 = x + 7

    x^2 - 3x - 40 = 0

    (x-8)*(x+5) = 0.


There are two roots x= 8  and x= -5.


Corresponding y-values are  y= 8+7 = 15  and  y= -5+7 = 2.


ANSWER.  The solutions are  (x,y) = (8,15)  and  (x,y) = (-5,2).

Solved.